40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5709     Entry time: Thu Oct 20 04:47:37 2011
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: clipping search round 1 

[Koji / Kiwamu]

  We tried finding a possible clipping in the vertex part.

We couldn't find an obvious location of a clipping but found that the recycling gain depended on the horizontal translation of the input beam.

We need more quantitative examination and should be able to find a sweet spot, where the recycling gain is maximized.

 

(what we did)

  + locked the carrier-resonant PRMI.

  + with IR viewers we looked at the inside of ITMX, ITMY and BS chambers to find an obvious clipping.

    => found two suspicious bright places and both were in the ITMY chamber.

      (1) POY pick off mirror : looked like a small portion of a beam was horizontally clipped by the mirror mount but not 100% sure whether if it is the main beam or a stray beam.

      (2) The top of an OSEM cable connectors tower : although this is in the way of the SRC path and nothing to do with PRC.

 + Made a hypothesis that the POY mirror is clipping the main beam.

 + To reject/prove the hypothesis we shifted the translation of the incident beam horizontally such that more beam hits on the suspicious mirror

 + Realigned and relocked PRMI.

    => Indeed the recycling gain went down from 6 to 0.8 or so. This number roughly corresponds to a loss of about 50%.

         However the MICH fringe still showed a very nice contrast (i.e. the dark fringe was still very dark).

         Therefore our conclusion is that the POY mirror is most likely innocent.

ELOG V3.1.3-